OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -4x - y greater than or equal to -8 A. Dashed line with shading on the side that includes the origin. B. Solid line with shading on the side that does not include the origin. C.Dashed line with shading on the side that does not include the origin. D. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 3x + 1 > 10 and 4x less than or greater to 20 A number line with an open circle on 3, a closed circle on 5, and shading in between. A number line with an open circle on 3, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on 3, shading to the left, and an open circle on 5, and shading to the right. A number line with a closed circle on 3, an open circle on 5, and shading in between.

OpenStudy (anonymous):

Help pleasee

OpenStudy (anonymous):

@zzr0ck3r ?? do u have answers?

OpenStudy (anonymous):

somoneone answer

Directrix (directrix):

For the first problem, because there is an "equal to" in the inequality statement, then the boundary will be "solid." Okay, that leaves B and D as possibilities. They differ only in whether or not the shaded section includes the origin. The origin has coordinates (0,0) Take this: -4x - y greater than or equal to -8 Replace both x and y by 0 and determine if the resulting inequality is true or false. Once you know that, you will know the correct option.

OpenStudy (anonymous):

Can you help me answer some more questions?

OpenStudy (anonymous):

oh and what about the second one

Directrix (directrix):

What did you get as an answer for the first question?

OpenStudy (anonymous):

i got B is that correct?

OpenStudy (anonymous):

it doesnt unclude origan

Directrix (directrix):

Why do you think that the origin is not to be included? What is the result of this: The origin has coordinates (0,0) Take this: -4x - y greater than or equal to -8 Replace both x and y by 0. Post what you get when you do that.

OpenStudy (anonymous):

-4(0) - 0>=-8 -4->=-8 It equals the same answer case watever times zero is that answer

Directrix (directrix):

I got this: -4x - y greater than or equal to -8 -4*0 - 0 >= to -8 ? 0 - 0 >= to -8 ? 0 >= -8 Yes

OpenStudy (anonymous):

so its B

Directrix (directrix):

All this -4*0 - 0 >= to -8 ? 0 - 0 >= to -8 ? 0 >= -8 True statement shows that the origin IS included in the shaded region because the coordinates of the origin make the inequality true. So, you want the option that INCLUDES the origin. Which one is it?

Directrix (directrix):

B is NOT correct.

OpenStudy (anonymous):

D includes the origan and its a solid line

Directrix (directrix):

Correct.

OpenStudy (anonymous):

What about the second question

OpenStudy (anonymous):

i need some help with others too if you dont mind